OpenStudy (anonymous):

how do I solve this system? 5x+4y=1 3x-4y=7

OpenStudy (anonymous):

add them together the 4y will cancel and your left with 8x = 8 or x=1 Then plug x = 1 into any of the equations and you get y = -1

OpenStudy (anonymous):

solve for why in both and set them equal to one another, y = 1/4 -5x/4 y= 3x/4 - 7/4 3x/4 - 7/4 = 1/4 - 5x/4 solve for x ==> x = 1 plug x back into either of the original equation y = 1/4 -5(1)/4 = -1

OpenStudy (anonymous):

Adding the equations is a bit nicer in this case: 5x + 4y = 1 + 3x - 4y = 7 =========== 8x + 0y = 8 \[\implies 8x = 8 \implies x = 1.\] Then plug in 1 to either original equation.